# On the Approach Towards Equilibrium Through Momentum-Dependent Relaxation: Insights from Evolution of the Moments in Kinetic Theory

### v2

@@ Abstract
We investigate the impact of momentum-dependent relaxation time approximation in the Boltzmann equation within the Bjorken flow framework by analyzing the moments of the single-particle distribution function. The moment equations, which form an infinite hierarchy, provide important insights about the system dynamics and the approach towards equilibrium for systems far from equilibrium. We show that a momentum-dependent collision kernel couples moments through both the energy exponents and the angular dependence via various-order Legendre polynomials, resulting in an intricate system of infinitely coupled equations that are complex and numerically challenging to solve. We outline strategies for solving the coupled system, including a novel approach for managing the infinite hierarchy and handling the non-integer moments. We show a significant influence of momentum dependent relaxation time on the time evolution of the moments, particularly for higher-order moments and system with smaller shear viscosity over entropy density, emphasizing the importance of incorporating such dependence for a more accurate description of the system dynamics with low shear viscosity such as the quark-gluon-plasma produced in high-energy heavy-ion collisions.
@@ I Introduction
Over the past few decades, relativistic hydrodynamics has found widespread applications in describing the dynamical evolution of collective systems, ranging from cosmology and astrophysics to the deconfined quark-gluon plasma produced in ultrarelativistic heavy-ion collisions [1, 2, 3]. This framework characterizes system evolution through conservation laws, leading to a set of coupled partial differential equations that govern macroscopic state variables and incorporate dissipative effects to account for irreversible phenomena.
@@
Despite the extensive success of relativistic hydrodynamics, the pathologies related to the physical properties have always troubled its journey. The first order relativistic Navier-Stokes (NS) theory [8, 9] being acausal with superluminal signal propagation and unstable against small perturbations[10, 11], the higher-order theories (mostly the Muller-Israel-Stewart (MIS) theory and its other variants [12, 14, 13, 15, 16]) are applied for the practical purposes like hydrodynamic simulations. However, they are applicable only within a strict range of its parameter space (in terms of its transport coefficients that can be derived from the underlying microscopic theory and act as input parameters of the hydro evolution equations) [17, 18]. Considering the scenario, a convenient alternative where these physical constraints can be bypassed could be to take suitable moments of the Boltzmann transport equation itself and recast it into an infinite hierarchy of differential equations of moments over the single-particle momentum distribution function. The method of moments has been known to be a useful technique for quite some time [19, 20, 21]. For a system where the moments do converge considerably so that the infinite hierarchy suffices to a limited number of the first few moments, this technique provides an alternative yet efficient way that saves from the complexities of the direct solution of the Boltzmann equation.
@@
In a number of recent works [22, 23, 24, 25], the method of moments has been studied to provide a quantitative measure of momentum anisotropies in a dissipative medium and to investigate the onset of hydrodynamics for a longitudinally expanding boost-invariant system described by the Bjorken flow [26]. In [27, 28] the convergence properties of the hierarchy of moment equations have been explored. In most of these studies, the source term in the kinetic equation (to be precise, the collision term of the relativistic Boltzmann equation) from which the moment hierarchy is built is expressed in terms of the relaxation time approximation (RTA) in its conventional form [29], that is, the relaxation time scale for the single-particle distribution function does not include the momentum dependence of the constituent particles.
@@
Now, the momentum dependence of the relaxation time is known to be related to microscopic interactions relevant to the medium under consideration [30]. Following this lead, in a number of recent studies [31, 32, 33, 34] the momentum-dependent relaxation time approximation has been adopted to solve the relativistic transport equation that captures the underlying momentum transfer to translate interesting features of the kinematics into the resulting hydrodynamic formulation. Considering the scenario, it is only indicative that momentum dependence should be included in the RTA approach for undertaking the moment analysis from the relativistic Boltzmann equation. The current analysis is an attempt in this direction where the sequence of moments has been derived from the Boltzmann equation using a momentum-dependent relaxation time approximation (MDRTA) with the boost invariant symmetry of Bjorken flow. The interesting features and the added complexities of this analysis are described below.
@@
In general (irrespective of the symmetry chosen) the moments are defined as a phase-space integral over the particle distribution function weighted by a product of a power of particle energy and an irreducible tensor of particle momenta having certain rank [27]. For a system that is invariant under Lorentz boosts described by Bjorken geometry [26] and consisting of massless particles (as considered in the present analysis), the definition of the moment $\rho_{n,l}$ takes the following form,
$\rho_{n,l}=\int\frac{d^{3}p}{(2\pi)^{3}p^{0}}\left(p^{0}\right)^{n}P_{2l}\left(\frac{p_{z}}{|\vec{p}|}\right)f_{p}\leavevmode\nobreak\ ,$ (1)
@@
with $f_{p}$ as the the single particle distribution function (with particle four-momenta $p^{\mu}$) and $P_{2l}$ is the Legendre polynomial with order $2l$. The equations of motion of $\rho_{n,l}$, which give the evolution dynamics for the moments derived from the kinetic theory, are observed to be highly coupled. In studies like [23, 27] it has been shown that the evolution equations of the lower moments duly depend on that of higher order, making a physical truncation of this moment hierarchy inevitable. In previous studies, it has also been observed that with momentum-independent RTA, only the indices $l$ of Eq.(1) are coupled. However, the energy exponent $n$ does not mix and attributes individual equations of motion for each moment. In the current analysis, the situation becomes much more involved when the particle momentum dependence is introduced in the relaxation time of the collision term in kinetic equation. The moments $\rho_{n,l}$ now get coupled via the $n$ indices as well through the momentum dependence of MDRTA. Consequently, the resulting chain of moment equations becomes much more interdependent and the attainment of convergence in such a scenario becomes much more complicated and numerically challenging. However, we argue that a momentum-dependent relaxation time approximation captures valuable dynamical information since (though via a simplistic model) it provides an account of the microscopic momentum transfer within the medium. The results corroborate our apprehension as we see that the inclusion of momentum dependence in the RTA formalism significantly impacts the behaviour of the moments as the time evolution of their solution becomes sensitive to the energy dependence of the relaxation time under MDRTA.
@@ II.1 The relativistic transport equation
In covariant kinetic theory, the microscopic behavior of a system of particles is characterized by the phase space distribution function $f_{p}(x^{\mu},p^{\mu})$, which, when multiplied with the appropriate phase space volume gives the probability of finding a particle at a point $(x^{\mu},p^{\mu})$ in phase space. The relativistic Boltzmann equation gives its evolution dynamics as the following,
$p^{\mu}\partial_{\mu}f_{p}(x^{\mu},p^{\mu})=C[f_{p}]\leavevmode\nobreak\ ,$ (2)
with $C[f_{p}]$ as the collision kernel that includes the system interactions. The Boltzmann equation is a non-linear integro-differential equation and therefore is both numerically and analytically challenging to solve. But to gain insight into the qualitative features of the solution we can approximate the collision kernel with model equations and convert it into a linearized operator as the following,
$p^{\mu}\partial_{\mu}f_{p}(x^{\mu},p^{\mu})={\cal{L}}[\phi_{p}]\leavevmode\nobreak\ .$ (3)
@@
$p^{\mu}\partial_{\mu}f_{p}(x^{\mu},p^{\mu})=-\frac{p^{\mu}u_{\mu}}{\tau_{R}}f_{p}^{eq}\leavevmode\nobreak\ \phi_{p}\leavevmode\nobreak\ .$ (4)
Here, the non-equilibrium distribution function $f_{p}$ is decomposed in an equilibrium and an out-of-equilibrium deviation part as $f_{p}-f_{p}^{eq}=f_{p}^{eq}\phi_{p}$, and ${\cal{L}}[\phi_{p}]$ is the linearized collision operator of this deviation function $\phi_{p}$. The equilibrium distribution function $f_{p}^{eq}$ is given by,
$f_{p}^{\text{eq}}=\frac{1}{e^{\left(p\cdot u-\mu\right)/T}+r}\leavevmode\nobreak\ ,$ (5)
where $r$ takes values $\{-1,0,1\}$ corresponding to Bose-Einstein, Boltzmann, and Fermi-Dirac statistics respectively. In the above definition, $u^{\mu}$ is a time like four vectors representing a local flow velocity, $T$ is the temperature and $\mu$ is the chemical potential of the system respectively.
Now, in conventional RTA formalism given by Anderson and Witting (AW-RTA) [29], the expression of $\tau_{R}$ in Eq.(4) is independent of particle momenta $p^{\mu}$. Hence, taking zeroth and first moment of Eq. (4) readily gives the conservation of particle-4-flow $N^{\mu}$ and energy-momentum tensor tensor $T^{\mu\nu}$ with the given form of the ${\cal{L}}[\phi_{p}]$ (right hand side of Eq.(4)) as long as the Landau matching condition of hydrodynamics is satisfied [4, 21]. Thus the AW-RTA Kernal is restricted to a specific choice of hydrodynamic frame. However, with a momentum dependent relaxation time approximation (MDRTA) (and with arbitrary hydrodynamic frames) the collision kernel given in Eq.(4) faces serious issues of conservation violation and thus unsuitable to study the non-equilibrium evolution of systems with MDRTA. The solution comes with the reformed collision kernels recently introduced and used in [33, 35, 36], that conserve the collision invariants irrespective of the momentum dependence in $\tau_{R}$ or the chosen hydrodynamic frame. The structural difference between the two referred collision kernels (under MDRTA) is merely because of the orthogonal basis in which the collision term is expanded. In the next subsection, when formulating the evolution dynamics of the moments, we will discuss this in more details.
@@ II.2.1 Basic definitions and properties
To study the evolution of bulk properties of the system, as mentioned in the introduction in Eq.(1), we define here the moments of the particle distribution function. But before that, we need to discuss the symmetries considered for the current analysis. In this study we consider a simplified conformal system having rotational and translational symmetry in the transverse $x-y$ plane and boost invariance along $z$ axis. These symmetries can be conveniently manifested using the hyperbolic (Milne) coordinates given by the metric tensor, $g_{\mu\nu}=(1,-1,-1,-\tau^{2})$ with proper time, $\tau=\sqrt{t^{2}-z^{2}}$ and space-time rapidity, $\eta_{s}=\tanh^{-1}{z/t}$. Under these constraints, the fluid flow profile reduces to $u^{\mu}=(1,0,0,0)$. The distribution function then only has spacetime dependence via $\tau$, and momentum dependence through the transverse momentum $p_{T}$ and longitudinal momentum $p_{z}=p_{\eta}/\tau$. Under these symmetry conditions the moments of particle distribution function takes the following form [27, 38, 39],
[…]
Here $dP=\frac{d^{3}p}{(2\pi)^{3}p^{0}}$ is the phase space factor, $E_{p}=p^{0}$ is the single particle energy and $\cos{\theta}=p_{z}/E_{p}=p_{z}/|\vec{p}|=p_{\eta}/(\tau p^{0})$ ). In the moment expression of (6) , the index $n$ measures energy scaling where the index $l$ measures momentum anisotropy in the system.
Next we list here few of the properties of the moment expression (6). The moment corresponding to equilibrium distribution function $\rho_{n,l}^{eq}$ has the following form,
@@
$\rho_{n,l}^{eq}(T,\mu)=e^{\mu/T}\frac{T^{n+2}}{2\pi^{2}}\Gamma(n+2)\delta_{l0}\leavevmode\nobreak\ ,$ (8)
@@
$\partial_{t}f_{p}+\vec{v_{p}}\cdot\vec{\nabla}f_{p}=\frac{1}{p^{0}}{\cal{L}}[\phi_{p}]\leavevmode\nobreak\ ,$ (9)
with $\vec{v}_{p}=\vec{p}/E_{p}$. Under Bjorken symmetry it takes the following form [37],
$\left[\frac{\partial}{\partial\tau}-\frac{1}{\tau}p_{z}\frac{\partial}{\partial p_{z}}\right]f_{p}=\frac{1}{p^{0}}{\cal{L}}[\phi_{p}]\leavevmode\nobreak\ .$ (10)

### v5

@@ Abstract
We investigate the impact of momentum-dependent relaxation time approximation in the Boltzmann equation within the Bjorken flow framework by analyzing the moments of the single-particle distribution function. The moment equations, which form an infinite hierarchy, provide important insights about the system dynamics and the approach towards equilibrium for systems far from equilibrium. The momentum-dependent collision kernel couples moments through both the energy exponents and the angular dependence via various-order Legendre polynomials, resulting in an intricate system of infinitely coupled equations. A naive truncation of the coupled equations results in diverging moments at late times. We outline strategies for solving the coupled system, including a novel approach for managing the divergences and non-integer moments. We show a significant influence of momentum dependent relaxation time on the time evolution of the moments, particularly for higher-order moments and system with smaller shear viscosity over entropy density, emphasizing the importance of incorporating such dependence for a more accurate description of the system dynamics with low shear viscosity such as the quark-gluon-plasma produced in high-energy heavy-ion collisions.
@@ I Introduction
Over the past few decades, relativistic hydrodynamics has found widespread applications in describing the dynamical evolution of collective systems, ranging from cosmology and astrophysics to the de-confined quark-gluon plasma produced in ultra-relativistic heavy-ion collisions [1, 2, 3]. This framework characterizes system evolution through conservation laws, leading to a set of coupled partial differential equations that govern macroscopic state variables and incorporate dissipative effects to account for irreversible phenomena.
@@
Despite the extensive success of relativistic hydrodynamics, pathologies related to its physical properties have always troubled its development. The first order relativistic Navier-Stokes (NS) theory [8, 9] is acausal with superluminal signal propagation and unstable against small perturbations[10, 11]. The higher-order theories, mostly the Muller-Israel-Stewart (MIS) theory and its other variants [12, 13, 14, 15, 16] partially solve these shortcomings and are applied for practical purposes like hydrodynamic simulations of heavy-ion collisions. However, they are applicable only within a strict range of its parameter space (in terms of its transport coefficients that can be derived from the underlying microscopic theory and act as input parameters of the hydro evolution equations) [17, 18]. Considering the scenario, a convenient alternative where these physical constraints can be bypassed could be to take suitable moments of the Boltzmann transport equation itself and recast it into an infinite hierarchy of differential equations of moments over the single-particle momentum distribution function. The method of moments has been known to be a useful technique for quite some time [19, 20, 21]. For a system where the moments converge fast enough, it suffices to consider a limited number of the first few moments in the infinite hierarchy. This technique provides an alternative yet efficient way that saves us the complexities of computing the exact solution of the Boltzmann equation.
@@
In several recent works [22, 23, 24, 25], the method of moments has been studied to provide a quantitative measure of momentum anisotropies in a dissipative medium and to investigate the onset of hydrodynamics for a longitudinally expanding boost-invariant system described by the Bjorken flow [26]. In [27, 28] the convergence properties of the hierarchy of moment equations have been explored. In most of these studies, the source term in the kinetic equation (to be precise, the collision term of the relativistic Boltzmann equation) is expressed in terms of the relaxation time approximation (RTA). In its conventional form [29], the relaxation time scale for the single-particle distribution function is taken to be independent of momenta.
@@
Now, the momentum dependence of the relaxation time is known to be related to microscopic interactions relevant to the medium under consideration [30]. Following this lead, in a number of recent studies [31, 32, 33, 34, 35, 36, 37, 38, 39, 40], the momentum-dependent relaxation time approximation has been adopted to derive the relativistic hydrodynamic equations. For example, in [31] Fig.[1] shows the dependence of differential scaled $v_{2}$ on the momentum-dependent correction to the distribution function. Considering the scenario, it is only indicative that momentum dependence should be included in the RTA approach for undertaking the moment analysis from the relativistic Boltzmann equation. The current analysis is an attempt in this direction where the sequence of moments has been derived from the Boltzmann equation using a momentum-dependent relaxation time approximation (MDRTA) with the boost invariant symmetry of Bjorken flow. The interesting features and the added complexities of this analysis are described below.
@@
In general (irrespective of the symmetry chosen) the moments are defined as a phase-space integral over the particle distribution function weighted by a product of a power of particle energy and an irreducible tensor of particle momenta having a certain rank [27]. For a system that is invariant under Lorentz boosts described by Bjorken geometry [26] and consisting of massless particles (as considered in the present analysis), the definition of the moment $\rho_{n,l}$ takes the following form,
$\rho_{n,l}=\int\frac{d^{3}p}{(2\pi)^{3}p^{0}}\left(p^{0}\right)^{n}P_{2l}\left(\frac{p_{z}}{|\vec{p}|}\right)f(x^{\mu},p^{\mu})\penalty 10000\ ,$ (1)
@@
with $f$ as the single particle distribution function (with particle four-momenta $p^{\mu}$) and $P_{2l}$ is the Legendre polynomial with order $2l$. The equations of motion of $\rho_{n,l}$, which give the evolution dynamics for the moments derived from the kinetic theory, are observed to be highly coupled. In studies like [23, 27] it has been shown that the evolution equations of the lower moments duly depend on that of higher order, making a physical truncation of this moment hierarchy inevitable. In previous studies, it has also been observed that with momentum-independent RTA, only the indices $l$ of Eq.(1) are coupled. However, the energy exponent $n$ does not mix and attributes individual equations of motion for each moment. In the current analysis, the situation becomes much more involved when the particle momentum dependence is introduced in the relaxation time of the collision term in kinetic equation. The moments $\rho_{n,l}$ now get coupled via the $n$ indices as well through the momentum dependence of MDRTA. Consequently, the resulting chain of moment equations becomes much more interdependent. However, we argue that a momentum-dependent relaxation time approximation captures valuable dynamical information since (though via a simplistic model) it provides an account of the microscopic momentum transfer within the medium. The results corroborate our apprehension as we see that the inclusion of momentum dependence in the RTA formalism significantly impacts the behavior of the moments as the time evolution of their solution becomes sensitive to the energy dependence of the relaxation time under MDRTA.
@@ II.1 The relativistic transport equation
In covariant kinetic theory, the microscopic behavior of a system of particles is characterized by the phase space distribution function $f(x^{\mu},p^{\mu})$, which, when multiplied with the appropriate phase space volume gives the probability of finding a particle at a point $(x^{\mu},p^{\mu})$ in phase space. The relativistic Boltzmann equation gives its evolution dynamics as the following,
$p^{\mu}\partial_{\mu}f_{p}(x^{\mu},p^{\mu})=C[f,f]\penalty 10000\ ,$ (2)
with $C[f,f]$ as the collision kernel that includes the system interactions. The Boltzmann equation is a non-linear integro-differential equation and therefore is both numerically and analytically challenging to solve. But to gain insight into the qualitative features of the solution we can approximate the collision kernel with model equations, the form of which is derived from the linearized operator,
$p^{\mu}\partial_{\mu}f(x^{\mu},p^{\mu})={\cal{L}}[\phi]\penalty 10000\ .$ (3)
@@
$p^{\mu}\partial_{\mu}f(x^{\mu},p^{\mu})=-\frac{p^{\mu}u_{\mu}}{\tau_{R}}f_{p}^{eq}\penalty 10000\ \phi\penalty 10000\ .$ (4)
Here, the non-equilibrium distribution function $f$ is decomposed in an equilibrium and an out-of-equilibrium deviation part as $f-f^{eq}=f^{eq}\phi$, and ${\cal{L}}[\phi]$ is the linearized collision operator of this deviation function $\phi$. The equilibrium distribution function $f^{eq}$ is given by,
$f^{\text{eq}}=\frac{1}{e^{\left(p\cdot u-\mu\right)/T}+r}\penalty 10000\ ,$ (5)
where $r$ takes values $\{-1,0,1\}$ corresponding to Bose-Einstein, Boltzmann, and Fermi-Dirac statistics, respectively. In the above definition, $u^{\mu}$ is a time like four vectors representing a local flow velocity, $T$ is the temperature and $\mu$ is the chemical potential of the system, respectively.
Now, in conventional RTA formalism given by Anderson and Witting (AW-RTA) [29], the expression of $\tau_{R}$ in Eq.(4) is independent of particle momenta $p^{\mu}$. Hence, taking zeroth and first moment of (4) readily gives the conservation of particle-4-flow $N^{\mu}$ and energy-momentum tensor tensor $T^{\mu\nu}$ with the given form of the ${\cal{L}}[\phi]$ (right hand side of Eq.(4)) as long as the Landau matching condition of hydrodynamics is satisfied [4, 21]. Thus the AW-RTA Kernel is restricted to a specific choice of hydrodynamic frame. However, with a momentum dependent relaxation time approximation (MDRTA) (and with arbitrary hydrodynamic frames) the collision kernel given in Eq.(4) faces serious issues of conservation violation and thus is unsuitable to study the non-equilibrium evolution of systems with MDRTA. The solution comes with the reformed collision kernels recently introduced and used in [33, 41, 42, 31], that conserve the collision invariants irrespective of the momentum dependence in $\tau_{R}$ or the chosen hydrodynamic frame. The structural difference between the two referred collision kernels (under MDRTA) is merely because of the orthogonal basis in which the collision term is expanded. In the next subsection, when formulating the evolution dynamics of the moments, we will discuss this in more details.
@@ II.2.1 Basic definitions and properties
To study the evolution of bulk properties of the system, as mentioned in the introduction in Eq.(1), we define here the moments of the particle distribution function. But before that, we need to discuss the symmetries considered for the current analysis. In this study we consider a simplified conformal system having rotational and translational symmetry in the transverse xyx-y plane and boost invariance along $z$ axis. These symmetries can be conveniently manifested using the hyperbolic (Milne) coordinates given by the metric tensor, $g_{\mu\nu}=(1,-1,-1,-\tau^{2})$ with proper time, $\tau=\sqrt{t^{2}-z^{2}}$ and space-time rapidity, $\eta_{s}=\tanh^{-1}{z/t}$. Under these constraints, the fluid flow profile reduces to $u^{\mu}=(1,0,0,0)$. The distribution function then only has spacetime dependence via $\tau$, and momentum dependence through the transverse momentum $p_{T}$ and longitudinal momentum $p_{z}=p_{\eta}/\tau$. Under these symmetry conditions, the moments of the particle distribution function take the following form [27, 43, 44],
@@
Next, we list here few of the properties of the moment expression (6). The moment corresponding to equilibrium distribution function $\rho_{n,l}^{eq}$ has the following form,
[…]
with the expression of $f_{p}^{eq}$ taken from (5) (we have used the Maxwell-Juttner distribution). Noticing that $f_{p}^{eq}$ is independent of $\theta$ (only function of $p$), we see that by the virtue of orthogonality property of Legendre polynomial $\int_{-1}^{1}P_{m}(x)P_{n}(x)dx\sim\delta_{mn}$, we have,
$\rho_{n,l}^{eq}(T,\mu)=e^{\mu/T}\frac{T^{n+2}}{2\pi^{2}}\Gamma(n+2)\delta_{l0}\penalty 10000\ ,$ (8)
@@
$\partial_{t}f_{p}+\vec{v_{p}}\cdot\vec{\nabla}f_{p}=\frac{1}{p^{0}}{\cal{L}}[\phi_{p}]\penalty 10000\ ,$ (9)
with $\vec{v}_{p}=\vec{p}/p$. Under Bjorken symmetry it takes the following form [45],
$\left[\frac{\partial}{\partial\tau}-\frac{p_{z}}{\tau}\frac{\partial}{\partial p_{z}}\right]f_{p}=-\frac{1}{p}{\cal{L}}[\phi_{p}]\penalty 10000\ .$ (10)